# Geometry of convex geometries

### v1

@@ 5 Realization of convex geometries and of their ideals
In this section we prove the main results of the paper. First, we prove that any convex geometry $\mathcal{A}=(U,\mathcal{C})$ admits a realization $(\mathcal{H}_{n},K(\mathcal{R}_{0}))$ in ${\mathbb{R}}^{n}$, where $n=|U|$, $\mathcal{H}_{n}$ are the coordinate hyperplanes of ${\mathbb{R}}^{n}$, and $K(\mathcal{R}_{0})$ has as many facets as the number of critical rooted circuits of $\mathcal{A}$. Then we prove that any ideal $\mathcal{I}$ of $\mathcal{A}$ is also realizable in ${\mathbb{R}}^{n}$ and the polyhedron $K_{0}$ realizing $\mathcal{I}$ has as many facets as the number of critical rooted circuits of $\mathcal{A}$ and positive circuits of $\mathcal{I}$. Furthermore, we prove that each convex geometry is realizable in dimension equal to its convex dimension. Finally, we prove that trees and multisimplicial complexes have realizations in dimension bounded linearly in their VC-dimension.
@@ 5.1 Realization of ideals of convex geometries
Let $U=\{1,\ldots,n\}$ and let $\mathcal{A}=(U,\mathcal{C})$ be a convex geometry on $U$ having $\mathcal{R}:=\mathcal{R}(\mathcal{A})$ as the set of rooted circuits and $\mathcal{R}_{0}:=\mathcal{R}_{0}(\mathcal{A})$ as the set of critical rooted circuits. For each rooted circuit $(C,r)$ of $\mathcal{A}$ consider the hyperplane $H(C,r)$ defined by the equation $|U|x_{r}=\sum_{e\in C\setminus\{r\}}x_{e}$. Denote by $K(C,r)$ the open halfspace of ${\mathbb{R}}^{n}$ determined by $H(C,r)$ as
$K(C,r)=\left\{x=(x_{1},\ldots,x_{d})\in{\mathbb{R}}^{n}:|U|x_{r}>\sum_{e\in C\setminus\{r\}}x_{e}\right\}.$
@@
$K(P)=\left\{x=(x_{1},\ldots,x_{d})\in{\mathbb{R}}^{d}:\sum_{e\in P}x_{e}<0\right\}.$
@@
We prove that the $X$-orthant $\mathcal{O}(X)$ intersects the convex set $K$, i.e., that there exists a point $x^{*}=(x^{*}_{1},\ldots,x^{*}_{n})\in\mathcal{O}(X)\cap K$. Let $|X|=k$ and suppose without loss of generality that $X=\{e_{1},\ldots,e_{k}\}$. First, we set $x^{*}_{e_{1}}=x^{*}_{e_{2}}=\cdots=x^{*}_{e_{k}}=1$. In the following, we set a negative value for each $x^{*}_{e},e\notin X$. Observe that for any rooted circuit $(C,r)$ such that $C\subseteq X$, we have $|U|x^{*}_{r}=|U|>(|C|-1)=\sum_{e\in C\setminus\{r\}}x_{e}$ and thus $x^{*}\in K(C,r)$, independently of the values of the remaining $n-k$ coordinates of $x^{*}$. Notice also that, since $X$ belongs to $\mathcal{C}^{\prime}$, $X$ does not contain any positive circuit of $\mathcal{C}^{\prime}$. Since $X\in\mathcal{C}$, by axiom (C3) of convex geometries, the elements of $U\setminus X$ can be ordered $e_{k+1},\ldots,e_{n}$, such that all the sets $X_{i}=X\cup\{x_{k+1},\ldots,x_{k+i}\}=X_{i-1}\cup\{x_{k+i}\},i=1,\ldots,n-k$ are convex sets of $\mathcal{C}$. We define the remaining $n-k$ coordinates of $x^{*}$ (which have to be negative) following the order $e_{k+1},\ldots,e_{n}$. Suppose that after $i-1$ steps the coordinates $x^{*}_{e_{k+1}},\ldots,x^{*}_{e_{k+i-1}}$ have been defined. At stage $i$ we need to define the value of $x^{*}_{e_{k+i}}$ in order to satisfy
  • all constraints of the form $|U|x^{*}_{r}>\sum_{e\in C\setminus\{r\}}x^{*}_{e}$, where $(C,r)\in\mathcal{R}(\mathcal{A})$, $e_{k+i}\in C$, and $C\subset X_{k+i}$;
@@
For any rooted circuit $(C,r)$ such that $e_{k+i}\in C\subseteq X_{k+i}$, the values of $x^{*}_{e}$ for all elements $e\in C\setminus\{e_{k+i}\}$ have been already defined. Since $X_{k+i-1}\in\mathcal{C}$ and since $(C,r)$ is a rooted circuit of $\mathcal{A}$, we know that $r\neq e_{k+i}$. Consequently, if we set $x^{*}_{e_{k+i}}<|U|x^{*}_{r}-\sum_{e\in C\setminus\{r,e_{k+i}\}}x^{*}_{e}$, the point $x^{*}$ is in the halfspace $K(C,r)$.
@@
Since $Y\notin\mathcal{C}$, by Theorem 13 there exists a rooted circuit $(C,r)$ such that $C\cap Y=C\setminus\{r\}$. Pick any point $y=(y_{1},\ldots,y_{n})$ from the $Y$-orthant $\mathcal{O}(Y)$ of ${\mathbb{R}}^{n}$ (i.e., $y_{e}>0$ for $e\in Y$ and $y_{e}<0$ for $e\notin Y$). Since $C\cap Y=C\setminus\{r\}$, we have $y_{r}<0$ and $y_{e}>0$ for any $e\in C\setminus\{r\}$. Consequently, $y$ does not satisfy the inequality $|U|y_{r}>\sum_{e\in C\setminus\{r\}}y_{e}$, and thus $y\notin K$. As a result, each orthant $\mathcal{O}(Y)$ with $Y\in\mathcal{C}^{*}$ is separated from $K(\mathcal{R})$ by a hyperplane of the form $H(C,r)$ with $(C,r)\in\mathcal{R}$. ∎
@@
For any $Y\in\mathcal{C}^{*}$, if there exists $r\in U\setminus Y$ such that $Y\cup\{r\}\in\mathcal{C}$, then there exists a critical rooted circuit $(C,r)$ such that $C\subseteq Y$, and thus for each $x\in K(\mathcal{R}_{0})$, we have $x_{r}>\sum_{g\in C\setminus\{r\}}\frac{1}{|U|}x_{g}$.
@@
Consider a minimal subset $Z\subseteq Y$ such that $Z\in\mathcal{C}^{*}$ and $Z\cup\{r\}\in\mathcal{C}$. Let $C=\operatorname{ex}(Z\cup\{r\})$ and note that $Z\cup\{r\}=\operatorname{conv}(C)$ by (C4). By the definition of $C$, for any $z\in C$, we have $Z\cup\{r\}\setminus\{z\}\in\mathcal{C}$, and by minimality of $Z$, we have $Z\setminus\{z\}\in\mathcal{C}$. Consequently, $(C,r)$ is a critical rooted circuit of $\mathcal{C}$ and thus by the definition of $K(\mathcal{R}_{0})$, for each $x\in K(\mathcal{R}_{0})$, we have $x_{r}>\sum_{g\in C\setminus\{r\}}\frac{1}{|U|}x_{g}$. ∎
@@
$=\left(\prod_{j=0}^{i+1}b_{j}\right)x_{e_{0}}+\sum_{g\in Y}\left(\sum_{j=0}^{i+1}\left(\prod_{\ell=j+1}^{i+1}b_{\ell}\right)a_{j,g}\right)x_{g}.\qed$
@@
$=\frac{1}{|Y|},\qed$
which concludes the proof of the lemma.
@@
The proof that $(\mathcal{H}_{n},K)$ is a realization of $\mathcal{I}$ is obtained by combining Lemmas 26, 27, and 28 and the equality $K=K(\mathcal{R})\cap K(\mathcal{P})$. The proof that $(\mathcal{H}_{n},K_{0})$ is a realization of $\mathcal{I}$ is obtained by combining Lemmas 26, 29, and 28 and the equality $K_{0}=K(\mathcal{R}_{0})\cap K(\mathcal{P})$. Finally, the proof of the first assertion of the theorem that the pairs $(\mathcal{H}_{n},K(\mathcal{R}))$ and $(\mathcal{H}_{n},K(\mathcal{R}_{0}))$ are realizations of $\mathcal{A}$ is a direct consequence of the second assertion because a convex geometry $\mathcal{A}$ is an ideal of itself not containing positive circuits (and thus $K=K(\mathcal{R})$ and $K_{0}=K(\mathcal{R}_{0})$). ∎
@@
$x_{1}+x_{3}$ $<3x_{2}$ $x_{2}+x_{4}$ $<3x_{3}$ $x_{1}+x_{4}$ $<3x_{2}$ $x_{1}+x_{4}$ $<3x_{3}$
On the other hand, among the rooted circuits of $\mathcal{A}$, only ($(\{1,2,3\},2)$ and ($(\{2,3,4\},3)$ are critical. The two first inequalities above correspond to these two critical rooted circuits. These two inequalities are sufficient to have a realization of $\mathcal{A}$ in $\mathbb{R}^{4}$. Indeed, from the two first inequalities, we obtain $3x_{1}+x_{4}<8x_{2}$ and $x_{1}+3x_{4}<8x_{3}$. This implies that there is no point in the realization corresponding to the sets $\{1,4\},\{1,2,4\}$, and $\{1,3,4\}$, that are precisely the subsets of $U$ that are forbidden by the circuits ($(\{1,2,4\},2)$, and ($(\{1,3,4\},3)$.
Figure 1: A convex geometry represented in black as subgraph of the 4-dimensional hypercube (in gray)
From Proposition 10, we obtain the following corollary.
@@
Any convex geometry $\mathcal{A}=(U,\mathcal{C})$ admits a point representation $\mathcal{M}(U,Q_{0})$ where $|Q_{0}|$ is the number of critical rooted circuits of $\mathcal{A}$.
@@ 5.2 Realization of convex geometries and convex dimension
Assume now that $C\in\mathcal{C}\setminus\{\varnothing\}$. For each $1\leq\ell\leq d$, let $e_{\ell}=\min_{\leq_{\ell}}C$. Observe that $U\setminus C=\bigcup_{\ell=1}^{d}\{e\in U:e\leq_{\ell}e_{\ell}\}$. Let $c_{\ell}=b_{\ell}(e_{\ell})$. If $e_{\ell}=\min_{\leq_{\ell}}U$, let $c^{-}_{\ell}=0$ and if $e_{\ell}\neq\min_{\leq_{\ell}}U$, let $e^{-}_{\ell}$ be the predecessor of $e_{\ell}$ in $\leq_{\ell}$ and let $c^{-}_{\ell}=b_{\ell}(e^{-}_{\ell})$. Consider the point $y\in K$ such that $y_{\ell}=1+c_{\ell}^{-}$ for each $1\leq\ell\leq d$. For each $e\in C$, note that $e_{\ell}\leq_{\ell}e$ for all $1\leq\ell\leq d$. Consequently, $c_{\ell}=b_{\ell}(e_{\ell})\leq b_{\ell}(e)$ for all $1\leq\ell\leq d$. Note that $c_{\ell}=d+1+dc_{\ell}^{-}=d(1+c_{\ell}^{-})+1>dy_{\ell}$. Consequently, $\sum_{\ell=1}^{d}\frac{y_{\ell}}{b_{\ell}(e)}\leq\sum_{\ell=1}^{d}\frac{y_{\ell}}{c_{\ell}}<\sum_{\ell=1}^{d}\frac{1}{d}=1$, establishing (i). For each $e\notin C$, there exists $\ell$ such that $e\leq_{\ell}e^{-}_{\ell}$. Therefore $b_{\ell}(e)\leq b_{\ell}(e^{-}_{\ell})=c_{\ell}^{-}<y_{\ell}$. Consequently, $\frac{y_{\ell}}{b_{\ell}(e)}>1$ and since $\frac{y_{i}}{b_{i}(e)}>0$ for all $1\leq i\leq d$ with $i\neq\ell$, we have $\sum_{i=1}^{d}\frac{y_{i}}{b_{i}(e)}>1$, establishing (ii). ∎
@@
Conversely, suppose that $e_{\ell}\leq_{\ell}e$ for all $1\leq\ell\leq d$. If there exists $\ell$ such that $e=e_{\ell}$, then by definition of $e_{\ell}$, we have $e\in C$. Assume now that for every $1\leq\ell\leq d$, $e_{\ell}\leq_{\ell}e$. For each $1\leq\ell\leq d$, let $e_{\ell}^{+}$ be the successor of $e_{\ell}$ in $\leq_{\ell}$ and note that $e_{\ell}^{+}\leq_{\ell}e$. Let $c_{\ell}=b_{\ell}(e_{\ell})$ and let $c_{\ell}^{+}=b_{\ell}(e_{\ell}^{+})$. By the definition of $e_{\ell}$, $e_{\ell}\in C$ and thus $\sum_{i=1}^{d}\frac{y_{i}}{b_{i}(e_{\ell})}<1$. Since $y\in K$, we have $\frac{y_{i}}{b_{i}(e_{\ell})}>0$ for each $1\leq i\leq d$ with $i\neq\ell$ and consequently, $\frac{y_{\ell}}{c_{\ell}}=\frac{y_{\ell}}{b_{\ell}(e_{\ell})}<1$. Observe that $c_{\ell}<c_{\ell}^{+}\leq b_{\ell}(e)$ for all $1\leq\ell\leq d$. Moreover, $c_{\ell}^{+}=dc_{\ell}+d+1$ and thus $b_{\ell}(e)\geq c_{\ell}^{+}>dc_{\ell}$. Consequently, $\sum_{\ell=1}^{d}\frac{y_{\ell}}{b_{\ell}(e)}<\sum_{\ell=1}^{d}\frac{y_{\ell}}{dc_{\ell}}=\frac{1}{d}\sum_{\ell=1}^{d}\frac{y_{\ell}}{c_{\ell}}<\frac{1}{d}d=1$. Therefore, $e\in C$ and thus we have $C=\{e\in U:e_{\ell}\leq_{\ell}e,\forall 1\leq\ell\leq d\}$, establishing that $C\in\mathcal{C}$. ∎
By Claim 35, for any convex set $C\in\mathcal{C}$, there exists a region corresponding to $C$ in $(K,\mathcal{H})$. By Claim 36, any region of $(K,\mathcal{H})$ corresponds to a convex $C$ of $\mathcal{A}=(U,\mathcal{C})$, concluding the proof of the theorem. ∎

### v2

@@ 5 Realization of convex geometries and of their ideals
In this section we prove the main results of the paper: any ideal $\mathcal{I}$ of $\mathcal{A}$ of a convex geometry $\mathcal{A}=(U,\mathcal{C})$ admits a realization $(\mathcal{H}_{n},K(\mathcal{R}_{0}))$ in ${\mathbb{R}}^{n}$, where $n=|U|$, $\mathcal{H}_{n}$ are the coordinate hyperplanes of ${\mathbb{R}}^{n}$, and the number of facets of $K(\mathcal{R}_{0})$ is at most the number of critical rooted circuits of $\mathcal{A}$ and positive circuits of $\mathcal{I}$. In particular, this yields a representation for convex geometries.
Furthermore, we prove that each convex geometry is realizable in dimension equal to its convex dimension. Finally, we prove that trees and multisimplicial complexes have realizations in dimension bounded linearly in their VC-dimension.
@@ 5.1 Realization of ideals of convex geometries
Let $U=\{1,\ldots,n\}$ and let $\mathcal{A}=(U,\mathcal{C})$ be a convex geometry on $U$ having $\mathcal{R}:=\mathcal{R}(\mathcal{A})$ as the set of rooted circuits and $\mathcal{R}_{0}:=\mathcal{R}_{0}(\mathcal{A})$ as the set of critical rooted circuits. For each rooted circuit $(C,r)$ of $\mathcal{A}$ consider the hyperplane $H(C,r)$ defined by the equation $nx_{r}=\sum_{e\in C\setminus\{r\}}x_{e}$. Denote by $K(C,r)$ the open halfspace of ${\mathbb{R}}^{n}$ determined by $H(C,r)$ as
$K(C,r)=\left\{x=(x_{1},\ldots,x_{n})\in{\mathbb{R}}^{n}:nx_{r}>\sum_{e\in C\setminus\{r\}}x_{e}\right\}.$
@@
$K(P)=\left\{x=(x_{1},\ldots,x_{n})\in{\mathbb{R}}^{n}:\sum_{e\in P}x_{e}<0\right\}.$
@@
We prove that the $X$-orthant $\mathcal{O}(X)$ intersects the convex set $K$, i.e., that there exists a point $x^{*}=(x^{*}_{1},\ldots,x^{*}_{n})\in\mathcal{O}(X)\cap K$. Let $|X|=k$ and suppose without loss of generality that $X=\{e_{1},\ldots,e_{k}\}$. First, we set $x^{*}_{e_{1}}=x^{*}_{e_{2}}=\cdots=x^{*}_{e_{k}}=1$. In the following, we set a negative value for each $x^{*}_{e},e\notin X$. Observe that for any rooted circuit $(C,r)$ such that $C\subseteq X$, we have $nx^{*}_{r}=n=|U|>(|C|-1)=\sum_{e\in C\setminus\{r\}}x_{e}$ and thus $x^{*}\in K(C,r)$, independently of the values of the remaining $n-k$ coordinates of $x^{*}$. Notice also that, since $X$ belongs to $\mathcal{C}^{\prime}$, $X$ does not contain any positive circuit of $\mathcal{C}^{\prime}$. Since $X\in\mathcal{C}$, by axiom (C3) of convex geometries, the elements of $U\setminus X$ can be ordered $e_{k+1},\ldots,e_{n}$, such that all the sets $X_{i}=X\cup\{x_{k+1},\ldots,x_{k+i}\}=X_{i-1}\cup\{x_{k+i}\},i=1,\ldots,n-k$ are convex sets of $\mathcal{C}$. We define the remaining $n-k$ coordinates of $x^{*}$ (which have to be negative) following the order $e_{k+1},\ldots,e_{n}$. Suppose that after $i-1$ steps the coordinates $x^{*}_{e_{k+1}},\ldots,x^{*}_{e_{k+i-1}}$ have been defined. At stage $i$ we need to define the value of $x^{*}_{e_{k+i}}$ in order to satisfy
  • all constraints of the form $nx^{*}_{r}>\sum_{e\in C\setminus\{r\}}x^{*}_{e}$, where $(C,r)\in\mathcal{R}(\mathcal{A})$, $e_{k+i}\in C$, and $C\subset X_{k+i}$;
@@
For any rooted circuit $(C,r)$ such that $e_{k+i}\in C\subseteq X_{k+i}$, the values of $x^{*}_{e}$ for all elements $e\in C\setminus\{e_{k+i}\}$ have been already defined. Since $X_{k+i-1}\in\mathcal{C}$ and since $(C,r)$ is a rooted circuit of $\mathcal{A}$, we know that $r\neq e_{k+i}$. Consequently, if we set $x^{*}_{e_{k+i}}<nx^{*}_{r}-\sum_{e\in C\setminus\{r,e_{k+i}\}}x^{*}_{e}$, the point $x^{*}$ is in the halfspace $K(C,r)$.
@@
Since $Y\notin\mathcal{C}$, by Theorem 15 there exists a rooted circuit $(C,r)$ such that $C\cap Y=C\setminus\{r\}$. Pick any point $y=(y_{1},\ldots,y_{n})$ from the $Y$-orthant $\mathcal{O}(Y)$ of ${\mathbb{R}}^{n}$ (i.e., $y_{e}>0$ for $e\in Y$ and $y_{e}<0$ for $e\notin Y$). Since $C\cap Y=C\setminus\{r\}$, we have $y_{r}<0$ and $y_{e}>0$ for any $e\in C\setminus\{r\}$. Consequently, $y$ does not satisfy the inequality $ny_{r}>\sum_{e\in C\setminus\{r\}}y_{e}$, and thus $y\notin K$. As a result, each orthant $\mathcal{O}(Y)$ with $Y\in\mathcal{C}^{*}$ is separated from $K(\mathcal{R})$ by a hyperplane of the form $H(C,r)$ with $(C,r)\in\mathcal{R}$. ∎
@@
For any $Y\in\mathcal{C}^{*}$, if there exists $r\in U\setminus Y$ such that $Y\cup\{r\}\in\mathcal{C}$, then there exists a critical rooted circuit $(C,r)$ such that $C\subseteq Y$, and thus for each $x\in K(\mathcal{R}_{0})$, we have $x_{r}>\sum_{g\in C\setminus\{r\}}\frac{1}{n}x_{g}$.
@@
Consider a minimal subset $Z\subseteq Y$ such that $Z\in\mathcal{C}^{*}$ and $Z\cup\{r\}\in\mathcal{C}$. Let $C=\operatorname{ex}(Z\cup\{r\})$ and note that $Z\cup\{r\}=\operatorname{conv}(C)$ by (C4). By the definition of $C$, for any $z\in C$, we have $Z\cup\{r\}\setminus\{z\}\in\mathcal{C}$, and by minimality of $Z$, we have $Z\setminus\{z\}\in\mathcal{C}$. Consequently, $(C,r)$ is a critical rooted circuit of $\mathcal{C}$ and thus by the definition of $K(\mathcal{R}_{0})$, for each $x\in K(\mathcal{R}_{0})$, we have $x_{r}>\sum_{g\in C\setminus\{r\}}\frac{1}{n}x_{g}$. $\Diamond$
@@
$=\left(\prod_{j=0}^{i+1}b_{j}\right)x_{e_{0}}+\sum_{g\in Y}\left(\sum_{j=0}^{i+1}\left(\prod_{\ell=j+1}^{i+1}b_{\ell}\right)a_{j,g}\right)x_{g}.$
$\Diamond$
@@
$=\frac{1}{|Y|},$
which concludes the proof of the lemma. ∎
@@
The proof that $(\mathcal{H}_{n},K)$ is a realization of $\mathcal{I}$ is obtained by combining Lemmas 27, 28, and 29 and the equality $K=K(\mathcal{R})\cap K(\mathcal{P})$. The proof that $(\mathcal{H}_{n},K_{0})$ is a realization of $\mathcal{I}$ is obtained by combining Lemmas 27, 30, and 29 and the equality $K_{0}=K(\mathcal{R}_{0})\cap K(\mathcal{P})$. Finally, the proof of the first assertion of the theorem that the pairs $(\mathcal{H}_{n},K(\mathcal{R}))$ and $(\mathcal{H}_{n},K(\mathcal{R}_{0}))$ are realizations of $\mathcal{A}$ is a direct consequence of the second assertion because a convex geometry $\mathcal{A}$ is an ideal of itself not containing positive circuits (and thus $K=K(\mathcal{R})$ and $K_{0}=K(\mathcal{R}_{0})$). ∎
@@
$x_{1}+x_{3}$ $<4x_{2}$ $x_{2}+x_{4}$ $<4x_{3}$ $x_{1}+x_{4}$ $<4x_{2}$ $x_{1}+x_{4}$ $<4x_{3}$
On the other hand, among the rooted circuits of $\mathcal{A}$, only ($(\{1,2,3\},2)$ and ($(\{2,3,4\},3)$ are critical. The two first inequalities above correspond to these two critical rooted circuits. These two inequalities are sufficient to have a realization of $\mathcal{A}$ in $\mathbb{R}^{4}$. Indeed, from the two first inequalities, we obtain $4x_{1}+x_{4}<15x_{2}$ and $x_{1}+4x_{4}<15x_{3}$. This implies that there is no point in the realization corresponding to the sets $\{1,4\},\{1,2,4\}$, and $\{1,3,4\}$, that are precisely the subsets of $U$ that are forbidden by the circuits ($(\{1,2,4\},2)$, and ($(\{1,3,4\},3)$.
From Proposition 12 and Proposition 14, we obtain the following corollary.
@@
Any convex geometry $\mathcal{A}=(U,\mathcal{C})$ admits a generalized convex shelling $\mathcal{C}(U,Q_{0})$ where $|Q_{0}|$ is the number of critical rooted circuits of $\mathcal{A}$.
@@ 5.2 Realization of convex geometries and convex dimension
Assume now that $C\in\mathcal{C}\setminus\{\varnothing\}$. For each $1\leq\ell\leq d$, let $e_{\ell}=\min_{\leq_{\ell}}C$. Observe that $U\setminus C=\bigcup_{\ell=1}^{d}\{e\in U:e\leq_{\ell}e_{\ell}\}$. Let $c_{\ell}=b_{\ell}(e_{\ell})$. If $e_{\ell}=\min_{\leq_{\ell}}U$, let $c^{-}_{\ell}=0$ and if $e_{\ell}\neq\min_{\leq_{\ell}}U$, let $e^{-}_{\ell}$ be the predecessor of $e_{\ell}$ in $\leq_{\ell}$ and let $c^{-}_{\ell}=b_{\ell}(e^{-}_{\ell})$. Consider the point $y\in K$ such that $y_{\ell}=1+c_{\ell}^{-}$ for each $1\leq\ell\leq d$. For each $e\in C$, note that $e_{\ell}\leq_{\ell}e$ for all $1\leq\ell\leq d$. Consequently, $c_{\ell}=b_{\ell}(e_{\ell})\leq b_{\ell}(e)$ for all $1\leq\ell\leq d$. Note that $c_{\ell}=d+1+dc_{\ell}^{-}=d(1+c_{\ell}^{-})+1>dy_{\ell}$. Consequently, $\sum_{\ell=1}^{d}\frac{y_{\ell}}{b_{\ell}(e)}\leq\sum_{\ell=1}^{d}\frac{y_{\ell}}{c_{\ell}}<\sum_{\ell=1}^{d}\frac{1}{d}=1$, establishing (i). For each $e\notin C$, there exists $\ell$ such that $e\leq_{\ell}e^{-}_{\ell}$. Therefore $b_{\ell}(e)\leq b_{\ell}(e^{-}_{\ell})=c_{\ell}^{-}<y_{\ell}$. Consequently, $\frac{y_{\ell}}{b_{\ell}(e)}>1$ and since $\frac{y_{i}}{b_{i}(e)}>0$ for all $1\leq i\leq d$ with $i\neq\ell$, we have $\sum_{i=1}^{d}\frac{y_{i}}{b_{i}(e)}>1$, establishing (ii). $\Diamond$
@@
Conversely, suppose that $e_{\ell}\leq_{\ell}e$ for all $1\leq\ell\leq d$. If there exists $\ell$ such that $e=e_{\ell}$, then by definition of $e_{\ell}$, we have $e\in C$. Assume now that for every $1\leq\ell\leq d$, $e_{\ell}\leq_{\ell}e$. For each $1\leq\ell\leq d$, let $e_{\ell}^{+}$ be the successor of $e_{\ell}$ in $\leq_{\ell}$ and note that $e_{\ell}^{+}\leq_{\ell}e$. Let $c_{\ell}=b_{\ell}(e_{\ell})$ and let $c_{\ell}^{+}=b_{\ell}(e_{\ell}^{+})$. By the definition of $e_{\ell}$, $e_{\ell}\in C$ and thus $\sum_{i=1}^{d}\frac{y_{i}}{b_{i}(e_{\ell})}<1$. Since $y\in K$, we have $\frac{y_{i}}{b_{i}(e_{\ell})}>0$ for each $1\leq i\leq d$ with $i\neq\ell$ and consequently, $\frac{y_{\ell}}{c_{\ell}}=\frac{y_{\ell}}{b_{\ell}(e_{\ell})}<1$. Observe that $c_{\ell}<c_{\ell}^{+}\leq b_{\ell}(e)$ for all $1\leq\ell\leq d$. Moreover, $c_{\ell}^{+}=dc_{\ell}+d+1$ and thus $b_{\ell}(e)\geq c_{\ell}^{+}>dc_{\ell}$. Consequently, $\sum_{\ell=1}^{d}\frac{y_{\ell}}{b_{\ell}(e)}<\sum_{\ell=1}^{d}\frac{y_{\ell}}{dc_{\ell}}=\frac{1}{d}\sum_{\ell=1}^{d}\frac{y_{\ell}}{c_{\ell}}<\frac{1}{d}d=1$. Therefore, $e\in C$ and thus we have $C=\{e\in U:e_{\ell}\leq_{\ell}e,\forall 1\leq\ell\leq d\}$, establishing that $C\in\mathcal{C}$. $\Diamond$
By 36, for any convex set $C\in\mathcal{C}$, there exists a region corresponding to $C$ in $(K,\mathcal{H})$. By 37, any region of $(K,\mathcal{H})$ corresponds to a convex $C$ of $\mathcal{A}=(U,\mathcal{C})$, concluding the proof of the theorem. ∎